# note-name = , use-sort-key = false

### v3

@@ III The Transient Phase-Space Dynamics
$\mathrm{d}_{\mathrm{tr}}\left(\rho(t),\rho^{\mathrm{(ss)}}\right)=\frac{1}{2}\mathrm{tr}\left({\left|\rho(t)-\rho^{\mathrm{(ss)}}\right|}\right)\leq\epsilon$ (14)
where $|\hat{A}|=\sqrt{\hat{A}^{\dagger}\hat{A}}$. The smallest $t$ for which this inequality is satisfied is what we call the steady-state time $T_{\mathrm{ss}}$. We choose the steady-state threshold $\epsilon$ to be small enough that $W$ and $P_{n}$ are sufficiently close to the steady-state values sy $T_{\mathrm{ss}}$, but not too small that the simulations take too long—choosing $\epsilon=10^{-3}$ achieves this.
@@ III.1 The Wigner function evolution
The system’s Wigner function evolves under the following equation of motion (see Appendix B for the derivation detail):
$\begin{split}\partial_{t}W&=\Bigg{\{}-\partial_{x}p+\partial_{p}x\\ &\qquad-\frac{1}{2}(\kappa_{1}-\gamma_{1})\left(\partial_{x}x+\partial_{p}p\right)+\frac{1}{4}(\kappa_{1}+\gamma_{1})\left(\partial_{x}^{2}+\partial_{p}^{2}\right)\\ &\qquad+\frac{\gamma}{2}\left[\partial_{x}\left(\left(x^{2}+p^{2}-2\right)x\right)+\partial_{p}\left(\left(x^{2}+p^{2}-2\right)p\right)+\left(\partial_{x}^{2}+\partial_{p}^{2}\right)\left(x^{2}+p^{2}-1\right)+\frac{1}{4}\left(\partial_{x}^{3}x+\partial_{x}^{2}\partial_{p}p+\partial_{x}\partial_{p}^{2}x+\partial_{p}^{3}p\right)\right]\Bigg{\}}W\end{split}$ (15)
The open system dynamics is stochastic in nature. It thus makes sense that the Wigner function evolves under an equation of motion that takes the form of the Kramers-Moyal (KM) equation [76, 77]. The KM equation describes the evolution of a probability distribution $P\left(\{\mathbb{x}\},t\right)$ of a state described by a set $\{\mathbb{x}\}=\{\mathbb{x}_{1},\mathbb{x}_{2},\dots,\mathbb{x}_{N}\}$ of Markovian stochastic variables or processes. It is given by
$\partial_{t}P\left(\{\mathbb{x}\},t\right)=\sum_{n=1}^{\infty}\partial_{\mathbb{x}_{j_{1}},\dots,\mathbb{x}_{j_{n}}}D^{(n)}_{\mathbb{x}_{j_{1}},\dots,\mathbb{x}_{j_{n}}}\left(\{\mathbb{x}\},t\right)P\left(\{\mathbb{x}\},t\right)$ (16)
@@
$D^{(n)}_{\mathbb{x}_{j_{1}},\dots,\mathbb{x}_{j_{n}}}\left(\{\mathbb{x}\},t\right)=\frac{(-1)^{n}}{n!}\lim_{\Delta t\rightarrow 0^{+}}\frac{1}{\Delta t}\left\langle{\mathcal{P}^{(n)}_{j_{1},\dots,j_{n}}\left(\Delta t;\{\mathbb{x}\},t\right)}\right\rangle$ (17)
@@
$\begin{split}\mathcal{P}^{(n)}_{j_{1},\dots,j_{n}}\left(\Delta t;\{\mathbb{x}\},t\right)&=\int_{-\infty}^{\infty}\mathrm{d}\xi_{j_{1}}\dots\mathrm{d}\xi_{j_{n}}\\ &\qquad\qquad\times\xi_{j_{1}}\dots\xi_{j_{n}}P\left(\{\mathbb{x}+\xi\},t+\Delta t\right)\end{split}$ (18)
@@
We divide the RHS of Eq. (15) into three rows. The first row is the simple harmonic part which consists of first-order derivatives, making them drift terms. Indeed, it is well known that the simple harmonic part is responsible for the rotation of $W$ about the phase-space origin [43].
@@
The second row takes the form of the Fokker-Planck equation, a special case of the KM equation where the RHS goes up to $n=2$ [76, 77]. The first-order derivatives make up the drift term, which shifts $W$ away from (if $\kappa_{1}>\gamma_{1}$) or toward (if $\kappa_{1}<\gamma_{1}$) the origin, corresponding to the exponential energy gain in Eq. (8). Meanwhile, the second-order derivatives make up the diffusion term, which increases the fuzziness of the system and is thus the embodiment of quantum decoherence imposed by the one-quantum processes. The rate of diffusion is constant.
@@
So far the equation of motion satisfies the properties of a KM equation. The third row, however, does not. The Pawula theorem states that the KM equation’s RHS ends at either $n=1$, $n=2$, or $n=\infty$ [77]. Furthermore, the coefficients $D^{(n)}_{\mathbb{x}_{1},\dots,\mathbb{x}_{n}}$ is nonnegative for even $n$ and $\mathbb{x}_{1}=\dots=\mathbb{x}_{n}$ as evident from Eqs. (17) and (18). The third row violates these properties. It ends with third-order derivatives and the second-order coefficient $x^{2}+p^{2}-1$ is not nonnegative. This deviation can be justified by noting that $W$ may take on negative values. This means that the propagator moment of $W$ need not be nonnegative for even $n$ and $\mathbb{x}_{1}=\dots=\mathbb{x}_{n}$ and that the Pawula theorem, which assumes that $P$ is nonnegative, does not apply here.
@@
We also discuss the continuity of the processes. As pointed out in Chapter 3 of Ref. [76], the KM equation’s RHS necessarily truncates at $n=2$ (the Fokker-Planck equation) for a continuous Markov processes, regardless of whether $P$ is nonnegative. This implies that the Markov process is jumpy if the KM equation truncates at $n>2$. In Chapter 4 of Ref. [76], it is shown that for nonnegative $P$ the KM equation never truncates for a jump Markov process—this, however, does not apply in our case so the previous statement is all we have. We can use these statements to interpret that the energy jumps imposed by the one-quantum processes correspond to small enough jumps in the quantum-phase-space variables that the processes are essentially continuous, resulting in $W$ evolving under a Fokker-Planck equation. Meanwhile, for higher-order energy jumps, the corresponding jumps in the quantum-phase-space variables are not small enough that the resulting contribution to the RHS of Eq. (15) truncates at $n>2$. In our case of two-quantum dissipation, it truncates at $n=3$.
@@
Figure 3: Wigner function $W(x,p)$ representation of the evolution of a coherent state $\left|{\beta}\right\rangle$ under different system parameters $(\kappa_{1},\gamma_{1},\gamma_{2},\beta)$: (a) $(1,0.1,0.05,5)$; (b) $(1,0.1,0.05,1)$; (c) $(1,0.9,0.005,5)$; (d) $(1,0.9,0.1,5)$; taken at different time points specified on top of each plot over the simple harmonic period of $2\pi$, alongside the trace distance $d_{\mathrm{tr}}$ toward the steady state given by Eq. (14). For each case, the time points are arbitrarily chosen to show: (1) the initial state, (2) some short time into the evolution before the SL phase point reaches its limit cycle, (3) the SL phase point reaching its limit cycle, (4) significant angular spreading of the Wigner function, and (5) when steady state is reached, as indicated by Eq. (14). For each plot, the color mapping is normalized against the Wigner function maximum at the given time point. The solid green lines show the Wigner function marginals, where the value 1 is set to be half the plot dimension. The dotted fuchsia circle is the classical SL limit cycle for the same parameters, given by $x^{2}+p^{2}=(\kappa_{1}-\gamma_{1})/\gamma_{2}$. The red dot gives the value of $\left\langle{a}\right\rangle$ obtained from $\rho(t)$ evolved under Eq. (1), while the dotted red line is its trail some time into the past. The orange square is the classical SL phase point $\alpha_{\mathrm{c}}(t)$ obtained by solving Eq. (4), while the dashed orange line is its trail some time into the past. Along the rightmost column are plots of the expected energy $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle$, and half the radial distance of the classical SL phase point from the origin which gives the energy of the classical SL oscillator, $E_{\mathrm{c}}=|\alpha_{\mathrm{c}}|^{2}$. The horizontal lines show the stationary values, while the vertical dotted black lines mark the time points of the Wigner function plots along the same row.
@@
We can see the requirement of nonlinearity for the emergence of the limit cycle in Eq. (15). For ease of visualization, we consider a vacuum $\rho=\left|{0}\right\rangle\left\langle{0}\right|$ whose Wigner function is a 2-dimensional Gaussian centered at the phase space origin. In this case, the second row only causes the Wigner function to be shifted radially outwards and spread radially uniformly with respect to its peak position at any given time. Without the third row, all we have is a Gaussian spiraling outward while getting spread out. It is the third row that enables nontrivial evolution features resulting in the limit cycle, such as localizing $W$ near the SL limit cycle amplitude. Unfortunately, we cannot tell which term does what due to the unconventional form of the third row.
@@ III.2 The classical-like behavior
It is a well-known property that under the simple harmonic Hamiltonian $\hat{H}_{0}=\hat{a}^{\dagger}\hat{a}+1/2$, the Wigner function rotates about the origin at the angular speed of $1\ \mathrm{rad/s}$. Specifically, for a coherent state $\left|{\beta}\right\rangle=\exp\left(-|\beta|^{2}/2\right)\sum_{n}\left(\beta^{n}/\sqrt{n!}\right)\left|{n}\right\rangle$, the Wigner function peak—which embodies the expected phase point $\left\langle{\hat{a}}\right\rangle=\left\langle{\beta}\middle|{\hat{a}}\middle|{\beta}\right\rangle=\beta$—traces out the classical circular trajectory [43]. Furthermore, its energy is given by $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle=\left\langle{\beta}\middle|{\hat{a}^{\dagger}\hat{a}}\middle|{\beta}\right\rangle=\left|\beta\right|^{2}$. The coherent state is the closest-to-classical quantum state, which makes it an appropriate choice to study the system’s behavior resembling the SL oscillator. We numerically simulate the evolution of the system’s Wigner function until it reaches the quantum limit cycle as indicated by Eq. (14), starting as a coherent state with different displacements. We also numerically simulate the SL oscillator’s phase point evolution initially at the peak position of the coherent state.
@@
Figure 3 presents the evolution snapshots in different working regimes and the corresponding energy evolution. We generally observe the Wigner function rotating at period $2\pi$ and a radial attraction similar to the SL oscillator. Regardless of the working regime, the radial attraction is similar in that the Wigner function’s peak coincides with the SL phase point. This exhibits some form of correspondence between the system and the SL oscillator even though the peak does not embody the quantity of interest, e.g. the energy. The coincidence can persist until the limit cycle in some regimes (not necessarily the classical regime, as shown in Fig. 2). In the other regimes, the coincidence is lost as the system evolves.
@@
The quantum decoherence caused by the nonconservative processes takes form as the angular spreading of the Wigner function resulting from the interplay between the nonconservative processes. Were it absent, we would expect the coherent state to keep its coherence and simply spiral toward the limit cycle amplitude. In reality, this angular spreading causes a gradual loss of information about the phase point, as shown by the expected phase point (red dot and dotted line) deviating from the SL trajectory (orange square and dashed line), and the Wigner function marginals (solid green plots)—which shows the position and momentum probability distribution—becoming broader and more symmetric over time. Ultimately, the steady state loses all the information about the phase point, being radially uniform. Nevertheless, the expected phase point agrees with the SL oscillator early in the evolution, regardless of the working regime, thanks to the initial coherence of the coherent state.
@@
By comparing the four cases shown in Fig. 3, we can see that evolutions in different working regimes are characterized by angular spreading of the Winger function. In the classical-eligible regime, the spreading along the radial direction is more confined as it approaches a steady-state distribution localized near the SL limit cycle. This also applies to cases where the system gains classical-regime eligibility somewhere along the evolution. As shown by case (b), the spreading is radially broader early in the evolution but becomes more localized after the system gains eligibility. On the contrary, in cases (c) and (d) the spreading is more localized early in the evolution but then becomes broader after the system loses eligibility.
@@
The energy evolution along the rightmost column of Fig. 3 is useful to see where the system loses eligibility and whether the evolution happens in the classical regime, as indicated by the energy evolution deviating from the SL oscillator. Case (a) shows agreement between the system’s and the SL oscillator’s energy evolution, characterizing that the system’s evolution is entirely in the classical regime. This result is interesting: if the system is eligible for all time, then starting as a coherent state, it entirely evolves in the classical regime. The angular spreading happens in such a way that the system’s energy keeps agreeing with the SL oscillator. In case (b), the system is initially not eligible and the energy evolution deviates from the SL oscillator, becoming similar after it gains eligibility and approaching the same limit cycle as case (a). In cases (c) and (d), the system’s energy only agrees with the SL oscillator early on, before deviating from it as the system loses eligibility.
@@
$\left(\Delta\rho(t)\right)_{mn}=\left|\left|\rho_{mn}(t)\right|-\left|\rho^{\mathrm{(ss)}}_{mn}\right|\right|$ (19)
For case (a) in Fig. 3, the result is shown in Fig. 4. The Cauchy-Schwarz inequality warrants that $\left|\rho_{mn}\right|^{2}\leq\rho_{mm}\rho_{nn}$—coherence only exists between two eigenstates with nonzero occupation probability. In our case, the coherence between high-energy eigenstates quickly vanishes. We see a swift “movement” toward lower-energy eigenstates, following $P_{n}$ as it quickly shifts due to the dominant nonlinear damping. Afterward, the coherence lingers in “the same region” as $P_{n}$ redistributes around $P_{n}^{\mathrm{(ss)}}$—here the energy gain is comparable to the dissipation hence no swift change in $P_{n}$. As $P_{n}$ settles down to $P_{n}^{\mathrm{(ss)}}$ in the fourth plot, the leftover coherence is between the eigenstates with nonzero occupation and other eigenstates that are a few energy levels away from them. From numerous simulations with different parameters and $\beta$ values—the result of which is provided in the Supplementary Materials—we find that the coherence between $\left|{n}\right\rangle$ and $\left|{n\pm 1}\right\rangle$ remains large compared to with higher-energy eigenstates, thus taking the longest to vanish. However, the initial magnitude of the $(n,n\pm 1)$ coherence is the largest, so this finding may not apply to an arbitrary quantum state.

### v5

@@ III Transient Phase-Space Dynamics
$\mathrm{d}_{\mathrm{tr}}\left(\rho(t),\rho^{\mathrm{(ss)}}\right)=\frac{1}{2}\mathrm{tr}\left({\left|\rho(t)-\rho^{\mathrm{(ss)}}\right|}\right)\leq\epsilon,$ (14)
where $|\hat{A}|=\sqrt{\hat{A}^{\dagger}\hat{A}}$. The smallest $t$ for which this inequality is satisfied is what we call steady-state time $T_{\mathrm{ss}}$. We choose the steady-state threshold $\epsilon$ to be small enough that $W$ and $P_{n}$ are sufficiently close to the steady-state values $T_{\mathrm{ss}}$, but not too small that the simulations take too long. Choosing $\epsilon=10^{-3}$ achieves this.
@@ III.1 The Wigner function evolution
The system’s Wigner function evolves under $\partial_{t}W=\hat{\mathcal{D}}_{\mathrm{(qSL)}}W$, where (see Appendix B for the derivation details):
$\begin{split}\hat{\mathcal{D}}_{\mathrm{(qSL)}}&=-\partial_{x}p+\partial_{p}x\\ &\qquad-\frac{1}{2}(\kappa_{1}-\gamma_{1})\left(\partial_{x}x+\partial_{p}p\right)+\frac{1}{4}(\kappa_{1}+\gamma_{1})\left(\partial_{x}^{2}+\partial_{p}^{2}\right)\\ &\qquad+\frac{\gamma_{2}}{2}\left\{\partial_{x}\left[\left(x^{2}+p^{2}-2\right)x\right]+\partial_{p}\left[\left(x^{2}+p^{2}-2\right)p\right]+\left(\partial_{x}^{2}+\partial_{p}^{2}\right)\left(x^{2}+p^{2}-1\right)+\frac{1}{4}\left(\partial_{x}^{3}x+\partial_{x}^{2}\partial_{p}p+\partial_{x}\partial_{p}^{2}x+\partial_{p}^{3}p\right)\right\}.\end{split}$ (15)
The open system dynamics is stochastic in nature. It thus makes sense that the Wigner function evolves under an equation of motion that resembles the Kramers-Moyal (KM) equation [77, 78]. The KM equation describes the evolution of a probability distribution $P\left(\{\mathbb{x}\},t\right)$ of a state described by a set $\{\mathbb{x}\}=\{\mathbb{x}_{1},\mathbb{x}_{2},\dots,\mathbb{x}_{N}\}$ of Markovian stochastic variables or processes. It is given by
$\partial_{t}P\left(\{\mathbb{x}\},t\right)=\sum_{n=1}^{\infty}\partial_{\mathbb{x}_{j_{1}},\dots,\mathbb{x}_{j_{n}}}D^{(n)}_{\mathbb{x}_{j_{1}},\dots,\mathbb{x}_{j_{n}}}\left(\{\mathbb{x}\},t\right)P\left(\{\mathbb{x}\},t\right),$ (16)
@@
$D^{(n)}_{\mathbb{x}_{j_{1}},\dots,\mathbb{x}_{j_{n}}}\left(\{\mathbb{x}\},t\right)=\frac{(-1)^{n}}{n!}\lim_{\Delta t\rightarrow 0^{+}}\frac{1}{\Delta t}\left\langle{\mathcal{P}^{(n)}_{j_{1},\dots,j_{n}}\left(\Delta t;\{\mathbb{x}\},t\right)}\right\rangle.$ (17)
@@
$\begin{split}\mathcal{P}^{(n)}_{j_{1},\dots,j_{n}}\left(\Delta t;\{\mathbb{x}\},t\right)&=\int_{-\infty}^{\infty}\mathrm{d}\xi_{j_{1}}\dots\mathrm{d}\xi_{j_{n}}\\ &\qquad\qquad\times\xi_{j_{1}}\dots\xi_{j_{n}}P\left(\{\mathbb{x}+\xi\},t+\Delta t\right),\end{split}$ (18)
@@
We divide the RHS of Eq. (15) into three rows. The first row represents the simple harmonic part, which consists of first-order derivatives, resulting in drift terms. Indeed, it is well known that the simple harmonic part is responsible for the rotation of $W$ about the phase-space origin [44].
@@
The second row takes the form of the Fokker-Planck equation, a special case of the KM equation where the RHS goes up to $n=2$ [77, 78]. The first-order derivatives make up the drift term, which shifts $W$ away from (if $\kappa_{1}>\gamma_{1}$) or toward (if $\kappa_{1}<\gamma_{1}$) the origin, corresponding to the exponential energy gain in Eq. (8). Meanwhile, the second-order derivatives comprise the constant-rate diffusion terms, which increase the fuzziness of the system and thus embody the quantum decoherence imposed by the one-quantum processes. This is well-known in the literature (see, e.g., Section 6.3.1 in Ref. [79]).
@@
So far, the equation of motion satisfies the properties of a KM equation. The third row, however, does not. The Pawula theorem states that the KM equation’s RHS ends at either $n=1$, $n=2$, or $n=\infty$ [78]. Furthermore, the coefficients $D^{(n)}_{\mathbb{x}_{1},\dots,\mathbb{x}_{n}}$ is nonnegative for even $n$ and $\mathbb{x}_{1}=\dots=\mathbb{x}_{n}$ as evident from Eqs. (17) and (18). The third row violates these properties. It ends with third-order derivatives, and the second-order coefficient $x^{2}+p^{2}-1$ is not nonnegative. This deviation can be justified by noting that $W$ may take on negative values. This means that the propagator moment of $W$ need not be nonnegative for even $n$ and $\mathbb{x}_{1}=\dots=\mathbb{x}_{n}$ and that the Pawula theorem, which assumes that $P\geq 0$, does not apply here.
@@
We also discuss the continuity of the processes. As pointed out in Chapter 3 of Ref. [77], the KM equation’s RHS necessarily truncates at $n=2$ (the Fokker-Planck equation) for a continuous Markov process, regardless of whether $P$ is nonnegative. This implies that the Markov process is jumpy if the KM equation truncates at $n>2$. In Chapter 4 of Ref. [77], it is shown that for nonnegative $P$ the KM equation never truncates for a jump Markov process. This, however, does not apply in our case, so the previous statement is all we have. We can use these statements to interpret that the energy jumps imposed by the one-quantum processes correspond to small enough jumps in the quantum-phase-space variables that the processes are essentially continuous, resulting in $W$ evolving under a Fokker-Planck equation. Meanwhile, for higher-order energy jumps, the corresponding jumps in the quantum-phase-space variables are not small enough that the resulting contribution to the RHS of Eq. (15) truncates at $n>2$. In our case of two-quantum dissipation, it truncates at $n=3$.
@@
Figure 3: Wigner function $W(x,p)$ representation of the evolution of a coherent state $\left|{\beta}\right\rangle$ under different system parameters $(\kappa_{1},\gamma_{1},\gamma_{2},\beta)$: (a) $(1,0.1,0.05,5)$; (b) $(1,0.1,0.05,1)$; (c) $(1,0.9,0.005,5)$; (d) $(1,0.9,0.1,5)$; taken at different time points specified on top of each plot over the simple harmonic period of $2\pi$, alongside the trace distance $d_{\mathrm{tr}}$ toward the steady state given by Eq. (14). For each case, the time points are arbitrarily chosen to show: (1) the initial state, (2) some short time into the evolution before the SL phase point reaches its limit cycle, (3) the SL phase point reaching its limit cycle, (4) significant angular spreading of the Wigner function, and (5) when steady state is reached, as indicated by Eq. (14). For each plot, the color mapping is normalized against the Wigner function maximum at the given time point. The solid green lines show the Wigner function marginals, where the value 1 is set to be half the plot dimension. The dotted fuchsia circle is the classical SL limit cycle for the same parameters, given by $x^{2}+p^{2}=(\kappa_{1}-\gamma_{1})/\gamma_{2}$. The red dot gives the value of $\left\langle{a}\right\rangle$ obtained from $\rho(t)$ evolved under Eq. (1), while the dotted red line is its trail some time into the past. The orange square is the classical SL phase point $\alpha_{\mathrm{c}}(t)$ obtained by solving Eq. (4), while the dashed orange line is its trail some time into the past. Along the rightmost column are plots of the expected energy $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle$, and half the radial distance of the classical SL phase point from the origin, which gives the energy of the classical SL oscillator, $E_{\mathrm{c}}=|\alpha_{\mathrm{c}}|^{2}$. The horizontal lines show the stationary values, while the vertical dotted black lines mark the time points of the Wigner function plots along the same row.
@@
We can see the requirement of nonlinearity for the emergence of the limit cycle in Eq. (15). For ease of visualization, we consider a vacuum $\rho=\left|{0}\right\rangle\left\langle{0}\right|$ whose Wigner function is a 2D Gaussian centered at the phase space origin. In this case, the second row only causes the Wigner function to be shifted radially outwards and spread radially uniformly with respect to its peak position at any given time. Without the third row, all we have is a Gaussian spiraling outward while getting spread out. It is the third row that enables non-trivial evolution features resulting in the limit cycle, such as localizing $W$ near the amplitude of the SL limit cycle. Unfortunately, we cannot identify the role of each term due to the unconventional form of the third row.
@@ III.2 The classical-like behavior
Under the simple harmonic Hamiltonian $\hat{H}_{0}=\hat{a}^{\dagger}\hat{a}+1/2$, the Wigner function is well-known to rotate about the origin at the angular speed of 1rad/s1\ $\mathrm${rad/s}. Specifically, for a coherent state $\left|{\beta}\right\rangle=\exp\left(-|\beta|^{2}/2\right)\sum_{n}\left(\beta^{n}/\sqrt{n!}\right)\left|{n}\right\rangle$, the Wigner function peak, which embodies the expected phase point $\left\langle{\hat{a}}\right\rangle=\left\langle{\beta}\middle|{\hat{a}}\middle|{\beta}\right\rangle=\beta$, traces out the classical circular trajectory [44]. Furthermore, its energy is given by $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle=\left\langle{\beta}\middle|{\hat{a}^{\dagger}\hat{a}}\middle|{\beta}\right\rangle=\left|\beta\right|^{2}$. The coherent state is the closest-to-classical quantum state, which makes it an appropriate choice to study the system’s behavior resembling the SL oscillator. We numerically simulate the evolution of the system’s Wigner function until it reaches the quantum limit cycle as indicated by Eq. (14), starting as a coherent state with different displacements. We also numerically simulate the evolution of the classical SL oscillator’s phase point, initialized at the peak position of the coherent state.
@@
Figure 3 presents the evolution snapshots in different working regimes and the corresponding energy evolution. We generally observe the Wigner function rotating at period $2\pi$ and a radial attraction similar to the SL oscillator. Regardless of the working regime, the radial attraction is similar in that the Wigner function’s peak coincides with the SL phase point. This exhibits some form of correspondence between the system and the SL oscillator, even though the peak does not embody the quantity of interest, e.g., the energy. The coincidence can persist until the limit cycle in some regimes (not necessarily the classical regime, as shown in Fig. 2). In the other regimes, the coincidence is lost as the system evolves.
@@
The quantum decoherence caused by the nonconservative processes takes the form of the angular spreading of the Wigner function resulting from the interplay between the nonconservative processes. Were it absent, we would expect the coherent state to keep its coherence and simply spiral toward the limit cycle amplitude. In reality, this angular spreading causes a gradual loss of information about the phase point, as shown by the expected phase point (red dot and dotted line) deviating from the SL trajectory (orange square and dashed line), and the Wigner function marginals (solid green plots), which shows the position and momentum probability distribution, becoming broader and more symmetric over time. Ultimately, the steady state loses all the information about the phase point, being radially uniform. Nevertheless, the expected phase point agrees with the SL oscillator early in the evolution, regardless of the working regime, thanks to the initial coherence of the coherent state.
@@
By comparing the four cases shown in Fig. 3, we can see that evolutions in different working regimes are characterized by the angular spreading of the Wigner function. In the classical-eligible regime, the spreading along the radial direction is more confined as it approaches a steady-state distribution localized near the SL limit cycle. This also applies to cases where the system gains classical-regime eligibility somewhere along the evolution. As shown by Fig. 3(b), the spreading is radially broader early in the evolution but becomes more localized after the system gains eligibility. On the contrary, in Figs. 3(c) and 3(d), the spreading is more localized early in the evolution but then becomes broader after the system loses eligibility.
@@
The energy evolution along the rightmost column of Fig. 3 is useful to see where the system loses eligibility and whether the evolution occurs in the classical regime, as indicated by the energy evolution deviating from the SL oscillator. The case of Fig. 3(a) shows agreement between the system’s and the SL oscillator’s energy evolution, characterizing that the system’s evolution is entirely in the classical regime. This result is interesting: if the system is eligible for all time, then starting as a coherent state, it entirely evolves in the classical regime. The angular spreading happens in such a way that the system’s energy keeps agreeing with the SL oscillator. In case (b), the system is initially not eligible, and the energy evolution deviates from the SL oscillator, becoming similar after it gains eligibility and approaching the same limit cycle as case (a). In cases (c) and (d), the system’s energy only agrees with the SL oscillator early on, before deviating from it as the system loses eligibility.
@@
[…]
For case (a) in Fig. 3, the result is shown in Fig. 4. The Cauchy-Schwarz inequality warrants that $\left|\rho_{mn}\right|^{2}\leq\rho_{mm}\rho_{nn}$, i.e., that coherence only exists between two eigenstates with nonzero occupation probability. In our case, the coherence between high-energy eigenstates quickly vanishes. We see a swift shift toward lower-energy eigenstates, following $P_{n}$ as it quickly shifts due to the dominant nonlinear damping. Afterward, the coherence lingers in the same region as $P_{n}$ redistributes around $P_{n}^{\mathrm{(ss)}}$. Here, the energy gain is comparable to the dissipation, hence no swift change in $P_{n}$. As $P_{n}$ settles down to $P_{n}^{\mathrm{(ss)}}$ in the fourth plot, the leftover coherence is between the eigenstates with nonzero occupation and other eigenstates that are a few energy levels away from them. From numerous simulations with different parameters and $\beta$ values, we find that the coherence between $\left|{n}\right\rangle$ and $\left|{n\pm 1}\right\rangle$ remains large compared to with higher-energy eigenstates, thus taking the longest to vanish. However, the initial magnitude of the $(n,n\pm 1)$ coherence is the largest, so this finding may not apply to an arbitrary quantum state.